E5-7 (Algo) Estimating Cost Behavior Using High-Low Method [LO 5-3]

Mountain Dental Services is a specialized dental practice whose only service is filling cavities. Mountain has recorded the following for the past nine months:

Month Number of Cavities Filled Total Cost
January 375 $5,300
February 500 5,850
March 350 5,200
April 600 6,250
May 325 5,150
June 475 5,700
July 525 6,100
August 575 6,300
September 450 5,550

Required:

1. Use the high-low method to estimate total fixed cost and variable cost per cavity filled.

2. Using these estimates, calculate Mountain’s total cost for filling 400 cavities.

Complete this question by entering your answers in the tabs below.

  • Required 1
  • Required 2

Use the high-low method to estimate total fixed cost and variable cost per cavity filled. (Round your Variable cost per unit to 2 decimal places.)

Fixed Cost

Variable Cost per cavity filled

                  

Using these estimates, calculate Mountain’s total cost for filling 400 cavities.

Estimated Total Cost                  
Answer #1

Variable cost = (Cost at highest activity - Cost at lowest activity) /(Highest activity-Lowest activity)

= (6250-5150)/(600-325) = 4

Total Cost = Variable + Fixed

6250 = (4*600) + Fixed

Fixed cost = 3850

Fixed cost 3850
Variable cost 4

Estimated total cost

= (400*4) + 3850

= 5450